OpenStudy (anonymous):

Evaluate the indicated sum of each of the arithmetic series given below. t1=11, t20=101, find S20

OpenStudy (nenadmatematika):

Sn=n(n+1)/2 so your sum is 20(20+1)/2=420/2=210

OpenStudy (nenadmatematika):

omg I'm wrong sorry here is the solution :D

OpenStudy (nenadmatematika):

Sn=(11+101)*(20/2)=112*10=1120

OpenStudy (nenadmatematika):

this is the formula for the sum of any arithmetic sequence: